OpenStudy (anonymous):

A wire carries a current of 5.2 A from east to west through a 6.0 m section of wire in the earth's magnetic field of 1 × 10^-5 T at a location where the magnetic field points south. What is the magnitude of the magnetic force acting on the wire? A. 3.1 × 10^-4 N B. 3.1 N C. 5.2 × 10^-5 N D. 6.0 × 10^-5 N

OpenStudy (michele_laino):

it is simple, since we have to apply this formula: \[\Large {\mathbf{F}} = I\;{\mathbf{l}} \times {\mathbf{B}}\] which can be rewritten as follows, since current and field are mutually perpendicular: \[\Large {\mathbf{F}} = I\;{\mathbf{l}} \times {\mathbf{B}} = IlB = 5.2 \times 6 \times {10^{ - 5}} = ...Newtons\]

OpenStudy (anonymous):

we get 3.12E-4? choice A is our solution?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay! thanks!!:) onto the next!:D

OpenStudy (michele_laino):

:) ok!